# Infinitary combinatorics in condensed mathematics and strong homology

### v1

@@ 2.3. The systems $\mathbf{A}_{\kappa,\lambda}$
Let $\kappa=|I|$ and $\lambda=|J|$ and observe that bijections $I\to\kappa$ and $J\to\lambda$ determine an isomorphism of the systems $\mathbf{A}_{I,J}$ and $\mathbf{A}_{\kappa,\lambda}$, so that we may without loss of generality restrict our attention to systems of the latter form. Thus for any nonzero cardinals $\kappa$ and $\lambda$ define inverse systems $\mathbf{B}_{\kappa,\lambda}=(B_{f},p_{fg},{{}^{\kappa}}([\lambda]^{<\omega})$ by letting
@@
and $\left(\mathbf{B}/\mathbf{A})_{\kappa,\lambda}=(B_{f}/A_{f},p_{fg},{{}^{\kappa}}([\lambda]^{<\omega}\right)$, where the maps $p_{fg}$ are in each case the obvious projection maps. Note that $\mathrm{lim}\,\mathbf{A}_{\kappa,\lambda}\cong\bigoplus_{\kappa}\prod_{\lambda}\mathbb{Z}$ and that $\mathrm{lim}\,\mathbf{B}_{\kappa,\lambda}\cong\prod_{\kappa\times\lambda}\mathbb{Z}$.
Write $\mathbf{X}\restriction\Lambda$ for the restriction of an inverse system $\mathbf{X}$ to a suborder $\Lambda$ of its index-set. Letting $[\omega]^{\mathrm{in}}$ denote the set of strictly initial segments of $\omega$ then induces an identification of $\mathbf{A}_{\omega,\omega}\restriction{{}^{\omega}}([\omega]^{\mathrm{in}})$ with the inverse system denoted $\mathbf{A}$ in [54] and [14], for example, and much intervening literature; since $[\omega]^{\mathrm{in}}$ is cofinal in $[\omega]^{<\omega}$, it follows that $\mathbf{A}_{\omega,\omega}$ and $\mathbf{A}$ are isomorphic in the category $\mathsf{pro}\text{-}\mathsf{Ab}$. The lemmas and definitions preceding Theorem 2.10 below are routine generalizations to the systems $\mathbf{A}_{\kappa,\lambda}$ of a sequence standard within the literature of $\mathbf{A}$.
@@
Observe that $\mathrm{lim}\,\mathbf{B}_{\kappa,\lambda}$ surjects onto $\mathrm{lim}\,(\mathbf{B}_{\kappa,\lambda}\restriction\Lambda)$ for any downwards-closed suborder $\Lambda$ of ([${{}^{\kappa}}([\lambda]^{<\omega})$; in other words, $\mathbf{B}_{\kappa,\lambda}$ is flasque in the sense of [46]. By way of this observation, the lemma follows from [46, Théorème 1.8]. ∎
@@
These equations facilitate combinatorial reformulations of the conditions “$\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}=0$” for $n>0$ as in the definition and lemma below; in what follows, $=^{*}$ denotes equality modulo a finite set.
@@
Let $\Phi=\langle\varphi_{f}:X(f)\to\mathbb{Z}\mid f\in{{}^{\kappa}}([\lambda]^{<\omega})\rangle$ be a family of functions.
@@
  • $\Phi$ is $n$-trivial if there exists a $\Psi=\langle\psi_{\vec{f}}:X(\vec{f})\to\mathbb{Z}\mid\vec{f}\in({{}^{\kappa}}([\lambda]^{<\omega}))^{n-1}\rangle$ such that
@@
The adaptations of these definitions to families of $H$-valued functions are straightforward, and the following lemma will continue to apply. For simplicity, however, we will tend in this subsection to proceed, as above, with a default codomain of $\mathbb{Z}$. Note also that coherence readily identifies in the above definition with the $n=1$ instance of $n$-coherence; this permits more uniform statements, as in the lemma below.
@@
For all $n>0$, $\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}=0$ if and only if every $n$-coherent family of functions indexed by ([${{}^{\kappa}}([\lambda]^{<\omega})$ is trivial.
@@
In the case of $n=1$, elements of $\mathrm{lim}\,(\mathbf{B}/\mathbf{A})_{\kappa,\lambda}$ may be viewed as collections of $=^{*}$-equivalence classes $\langle[\varphi_{f}]\mid f\in{{}^{\kappa}}([\lambda]^{<\omega})\rangle$, where $\langle\varphi_{f}\mid f\in{{}^{\kappa}}([\lambda]^{<\omega})\rangle$ is a coherent family of functions. Since $\mathrm{lim}\,\mathbf{B}_{\kappa,\lambda}=\prod_{\kappa\times\lambda}\mathbb{Z}$ and the map $\mathrm{lim}\,\mathbf{B}_{\kappa,\lambda}\to\mathrm{lim}\,(\mathbf{B}/\mathbf{A})_{\kappa,\lambda}$ takes a function $\psi:\kappa\times\lambda\to\mathbb{Z}$ to the collection $\langle[\psi|_{X(f)}]\mid f\in{{}^{\kappa}}([\lambda]^{<\omega})\rangle$ of equivalence classes of its restrictions, the assertion follows from equation 6.
The case of higher $n$ is close in spirit, only a bit more tedious; since the argument is so essentially that given in [16, Section 2.1], readers are referred there for details. The fundamental point is that, by equation 7, elements of $\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}$ admit representation as elements of $\mathrm{lim}^{n-1}\,(\mathbf{B}/\mathbf{A})_{\kappa,\lambda}$ which, when construed as cohomology classes in the associated Roos complex, give rise to the representations of Definition 2.8. ∎
@@
Item (1) is immediate from Definition 2.8 and Lemma 2.9, for if $\kappa$ is finite then any associated $n$-coherent function is trivialized by a family of zero functions (or by a single zero function, if $n=1$), while if $\lambda$ is finite then the partial order ([${{}^{\kappa}}([\lambda]^{<\omega})$ possesses a maximum.
@@
For item (4), fix $n\geq 0$ and observe that for any $\mu\geq\kappa$ and $\nu\geq\lambda$, any $n$-coherent $\Phi=\langle\varphi_{\vec{f}}:X(\vec{f})\to\mathbb{Z}\mid\vec{f}\in{{}^{\kappa}}([\lambda]^{<\omega})\rangle$ naturally extends to an $n$-coherent $\Psi=\langle\psi_{\vec{f}}:X(\vec{f})\to\mathbb{Z}\mid\vec{f}\in{{}^{\mu}}([\nu]^{<\omega})\rangle$ which is trivial if and only if $\Phi$ is, an extension which determines, in turn, an embedding of $\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}$ into $\mathrm{lim}^{n}\,\mathbf{A}_{\mu,\nu}$. Concretely, for any $f:\mu\to[\nu]^{<\omega}$ define $g:\kappa\to[\lambda]^{<\omega}$ by $g(\xi)=f(\xi)\cap\lambda$ for all $\xi\in\kappa$, a correspondence extending in an obvious fashion to $n$-tuples $\vec{f}\in({{}^{\mu}}([\nu]^{<\omega}))^{n}$. Let then $\psi_{\vec{f}}\,|_{X(\vec{g})}=\varphi_{\vec{g}}$ and $\psi_{\vec{f}}\,|_{X(\vec{f})\backslash X(\vec{g})}=0$; the verification that $\Psi$ is as claimed is straightforward and left to the reader.
@@
for all $(i,\xi)\in X(f)$. To complete the proof, it will suffice to show that $\Phi=\langle\varphi_{f}\mid f\in{{}^{\omega}}([\omega_{1}]^{<\omega})\rangle$ is nontrivially coherent. To aid in seeing this, consider an auxiliary family $T=\langle\tau_{\gamma}\mid\gamma<\omega_{1}\rangle$ of functions $\tau_{\gamma}:\omega\times\gamma\to\{0,1\}$, each of which is the characteristic function of the reflection of the graph of $e_{\gamma}$; more concisely, $\tau_{\gamma}(i,\xi)=1$ if and only if $e_{\gamma}(\xi)=i$. It is then clear from the observation that $\varphi_{f}=\tau_{\mathrm{sp}(f)}|_{X(f)}$ that $\Phi$, by way of $T$, inherits the coherence of $E$.
@@ 2.4. An alternative framing
    Whenever $\mathbf{M}=(M_{i},\pi_{i,j},\omega)$ is an inverse sequence $\mathbf{M}=(M_{i},\pi_{i,j},\omega)$ of abelian groups of cardinality at most $\lambda$ whose transition maps $\pi_{i,j}:M_{j}\to M_{i}$ are all surjective, and $H$ is an abelian group of cardinality at most $\mu$, we have
@@
of each $M_{i}$ in which each $F_{i}$ and $K_{i}$ are free groups of size $\lambda$ (i.e., each $F_{i}$ and $K_{i}$ are of the form $\bigoplus_{\lambda}\mathbb{Z}$) and observe that since products are exact in $\mathsf{Cond(Ab)}$ ([74, Thm. 2.2 (AB4*)]), these resolutions assemble into a short exact sequence
@@
The first is an easy consequence of the theorem’s clause 2, while the second further leverages the slenderness of $\bigoplus_{\mu}\mathbb{Z}$ [36, §94]. For the aforementioned projection maps from (9) to the condensations of (8) induce a chain map from the sum of the long exact sequences given by $\mathrm{Hom}(-,\bigoplus_{\mu}\underline{\mathbb{Z}})$ of the latter to that given by $\mathrm{Hom}(-,\bigoplus_{\mu}\underline{\mathbb{Z}})$ of the former, and slenderness tells us that the two constituents
@@
00{0}${\lim_{i<\omega}\,\underline{M_{i}}}$${\prod_{i<\omega}\underline{M_{i}}}$${\prod_{i<\omega}\underline{M_{i}}}$00{0}00{0}${\underline{M_{j}}}$${\prod_{i<j+1}\underline{M_{i}}}$${\prod_{i<j}\underline{M_{i}}}$00{0}$\scriptstyle{\mathrm{id}-\mathrm{sh}}$$\scriptstyle{\mathrm{id}-\mathrm{sh}}$
@@
Furthermore, a result of Casarosa and Lambie-Hanson [21] shows that, again for $\lambda=\aleph_{0}$, $\aleph_{\omega+1}$ is the minimum value of the continuum compatible with the conditions holding simultaneously for all $\mu$. They prove that, if the dominating number $\mathfrak{d}$ equals $\aleph_{n}$ for some $1\leq n<\omega$, then $\lim^{n}\mathbf{A}_{\aleph_{0},\aleph_{0}}[\bigoplus_{\omega_{n}}\mathbb{Z}]\neq 0$. In particular, if the conditions in Theorem 2.13 hold for $\lambda=\aleph_{0}$ and all cardinals $\mu<\aleph_{\omega}$, then we must have $2^{\aleph_{0}}>\aleph_{\omega}$. A more axiomatic approach to these results is possible as well; the assumption $[\mathrm{PH}_{n}$ for all $n\geq 0]$, where $\mathrm{PH}_{n}$ denotes the $n^{\mathrm{th}}$ partition hypothesis of [7], also implies Theorem 2.13 for $\lambda=\aleph_{0}$ and all $\mu>0$ simultaneously. This, though, is a logically stronger assumption than those discussed above: by [7, Cor. 8.21], its consistency strength is exactly that of the existence of a weakly compact cardinal.
@@
We end this subsection by highlighting a particular consequence of the assertions in Theorem 2.13 for the duality of condensed Banach spaces and condensed Smith spaces in non-archimedean settings. For concreteness, we work over $\mathbb{Q}_{p}$ for some fixed prime $p$. As noted in [25, Lecture 6], the derived category $\mathsf{D(Solid}_{\mathbb{Q}_{p}})$ of solid $\mathbb{Q}_{p}$-vector spaces is a full subcategory of the derived category $\mathsf{D(Solid)}$ of solid abelian groups, which in turn is a full subcategory of the the derived category $\mathsf{D(Cond}(\mathsf{Ab}))$ of condensed abelian groups.
@@
As shown in [71, Lemma 3.8] (see also [25, Lecture 6], which more closely matches our notation), the solid $\mathbb{Q}_{p}$-Banach spaces are precisely the solid $\mathbb{Q}_{p}$-vector spaces of the form
@@
for some index set $I$ (readability dictates that for the remainder of this section we forego the underline notation for condensed images of rings). It is shown in [71, Lemma 3.10] that the classical Banach–Smith duality extends to the solid $\mathbb{Q}_{p}$-vector space setting via the internal $\mathrm{Hom}$ functor $\underline{\mathrm{Hom}}(-,-):(\mathsf{Solid}_{\mathbb{Q}_{p}})^{\mathrm{op}}\times\mathsf{Solid}_{\mathbb{Q}_{p}}\rightarrow\mathsf{Solid}_{\mathbb{Q}_{p}}$ (see [74, p. 13] for basic definitions of condensed internal $\mathrm{Hom}$ functors). In particular, there it is shown that, for any index set $I$, we have
@@
As $\prod_{I}\mathbb{Z}_{p}$ is, moreover, a projective solid $\mathbb{Z}_{p}$-module, it is essentially immediate that the second of these equalities holds even at the level of the derived category $\mathsf{D(Solid}_{\mathbb{Q}_{p}})$, i.e., even with the $\underline{\mathrm{Hom}}$ replaced by $\mathrm{R}\underline{\mathrm{Hom}}$; see [71, Remark 3.11]. The question of whether this holds for the first bullet (and hence of whether this is a derived duality relation), though, is more delicate. We end this subsection with a brief proof of the observation of Clausen and Scholze (again, see [25, Lecture 6]) that, under the conditions of Theorem 2.13 with $\lambda=\aleph_{0}$, this question also has an affirmative answer in the context of separable solid $\mathbb{Q}_{p}$-Banach spaces, i.e., in the situation in which the index set $I$ is countable. The following theorem yields clause (3) of Theorem B:
@@
By [71, Lemma 3.10], we have $\underline{\mathrm{Hom}}\left((\bigoplus_{\omega}\mathbb{Z}_{p})^{\wedge}_{p}\left[\frac{1}{p}\right],\mathbb{Q}_{p}\right)=(\prod_{\omega}\mathbb{Z}_{p})\left[\frac{1}{p}\right]$. Therefore, it suffices to show that $\mathrm{R}\underline{\mathrm{Hom}}\left((\bigoplus_{\omega}\mathbb{Z}_{p})^{\wedge}_{p}\left[\frac{1}{p}\right],\mathbb{Q}_{p}\right)$ is concentrated in degree zero. Note that
@@
For the second equality, see the third paragraph of [79, §4] (or, a little more concretely, argue as for [17, Lem. 11.1]). Since $S\in\mathsf{ED}$ and $\mathbb{Q}_{p}/(p^{m}\mathbb{Z}_{p})$ is discrete, we know that $\mathrm{Cont}(S,\mathbb{Q}_{p}/(p^{m}\mathbb{Z}_{p}))$ (with the compact-open topology) is a discrete abelian group for all $m<\omega$. Hence clause (3) of Theorem 2.13 applies to the interior RHom of line 2.4, entailing that
$H^{i}\mathrm{RHom}\left(\lim_{n<\omega}\Big{(}\bigoplus_{\omega}\mathbb{Q}_{p}/(p^{n}\mathbb{Z}_{p})\Big{)},\mathrm{Cont}(S,\mathbb{Q}_{p}/(p^{m}\mathbb{Z}_{p}))\right)$

### v2

@@ 2.3. The systems $\mathbf{A}_{\kappa,\lambda}$
Let $\kappa=|I|$ and $\lambda=|J|$ and observe that bijections $I\to\kappa$ and $J\to\lambda$ determine an isomorphism of the systems $\mathbf{A}_{I,J}$ and $\mathbf{A}_{\kappa,\lambda}$, so that we may without loss of generality restrict our attention to systems of the latter form. Thus for any nonzero cardinals $\kappa$ and $\lambda$ define inverse systems $)\mathbf{B}_{\kappa,\lambda}=(B_{f},p_{fg},{{}^{\kappa}}([\lambda]^{<\omega}$)) by letting
@@
and $)\left(\mathbf{B}/\mathbf{A})_{\kappa,\lambda}=(B_{f}/A_{f},p_{fg},{{}^{\kappa}}([\lambda]^{<\omega}\right$)), where the maps $p_{fg}$ are in each case the obvious projection maps. Note that $\mathrm{lim}\,\mathbf{A}_{\kappa,\lambda}\cong\bigoplus_{\kappa}\prod_{\lambda}\mathbb{Z}$ and that $\mathrm{lim}\,\mathbf{B}_{\kappa,\lambda}\cong\prod_{\kappa\times\lambda}\mathbb{Z}$.
[…]
@@
[…]
@@
These equations facilitate combinatorial reformulations of the conditions
$\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}=0$"
for $n>0$ as in the definition and lemma below; in what follows, $^{*}$ denotes equality modulo a finite set.
@@
[…]
@@
  • $\Phi$ is $n$-trivial (or simply trivial if the value of $n$ is clear from context) if there exists a $\Psi=\langle\psi_{\vec{f}}:X(\vec{f})\to\mathbb{Z}\mid\vec{f}\in({{}^{\kappa}}([\lambda]^{<\omega}))^{n-1}\rangle$ such that
@@
The adaptations of these definitions to families of $H$-valued functions for an arbitrary $H\in\mathsf{Ab}$ are straightforward, and the following lemma will continue to apply. For simplicity, however, we will tend in this subsection to proceed, as above, with a default codomain of $\mathbb{Z}$. Note also that coherence readily identifies in the above definition with the $n=1$ instance of $n$-coherence; this permits more uniform statements, as in the lemma below.
@@
[…]
@@
[…]
The case of higher $n$ is close in spirit, only a bit more tedious; since the argument is so essentially that given in [18, Section 2.1], readers are referred there for details. The fundamental point is that, by equation 7, elements of $\mathrm{lim}^{n}\,\mathbf{A}_{\kappa,\lambda}$ admit representation as elements of $\mathrm{lim}^{n-1}\,(\mathbf{B}/\mathbf{A})_{\kappa,\lambda}$ which, when construed as cohomology classes in the associated Roos complex, give rise to the representations of Definition 2.8. ∎
@@
[…]
@@
[…]
@@
[…]
@@ 2.4. An alternative framing
    Whenever $\mathbf{M}=(M_{i},\pi_{i,j},\omega)$ is an inverse sequence of abelian groups of cardinality at most $\lambda$ whose transition maps $\pi_{i,j}:M_{j}\to M_{i}$ are all surjective, and $H$ is an abelian group of cardinality at most $\mu$, we have
@@
of each $M_{i}$ in which each $F_{i}$ and $K_{i}$ are free groups of size $\lambda$ (i.e., each $F_{i}$ and $K_{i}$ are of the form $\bigoplus_{\lambda}\mathbb{Z}$) and observe that since products are exact in $\mathsf{Cond(Ab)}$ ([77, Thm. 2.2 (AB4*)]), these resolutions assemble into a short exact sequence
@@
The first is an easy consequence of the theorem’s clause 2, while the second further leverages the slenderness of $\bigoplus_{\mu}\mathbb{Z}$ [39, §94]. For the aforementioned projection maps from (9) to the condensations of (8) induce a chain map from the sum of the long exact sequences given by $\mathrm{Hom}(-,\bigoplus_{\mu}\underline{\mathbb{Z}})$ of the latter to that given by $\mathrm{Hom}(-,\bigoplus_{\mu}\underline{\mathbb{Z}})$ of the former, and slenderness tells us that the two constituents
@@
[…]
@@
Furthermore, a result of Casarosa and Lambie-Hanson [23] shows that, again for $\lambda=\aleph_{0}$, $\aleph_{\omega+1}$ is the minimum value of the continuum compatible with the conditions holding simultaneously for all $\mu$. They prove that, if the dominating number $\mathfrak{d}$ equals $\aleph_{n}$ for some $1\leq n<\omega$, then $\lim^{n}\mathbf{A}_{\aleph_{0},\aleph_{0}}[\bigoplus_{\omega_{n}}\mathbb{Z}]\neq 0$. In particular, if the conditions in Theorem 2.13 hold for $\lambda=\aleph_{0}$ and all cardinals $\mu<\aleph_{\omega}$, then we must have $2^{\aleph_{0}}>\aleph_{\omega}$. A more axiomatic approach to these results is possible as well; the assumption $\mathrm{PH}_{n}$ for all $n\geq 0]$, where $\mathrm{PH}_{n}$ denotes the $n^{\mathrm{th}}$ partition hypothesis of [8], also implies Theorem 2.13 for $\lambda=\aleph_{0}$ and all $\mu>0$ simultaneously. This, though, is a logically stronger assumption than those discussed above: by [8, Cor. 8.21], its consistency strength is exactly that of the existence of a weakly compact cardinal.
@@
We end this subsection by highlighting a particular consequence of the assertions in Theorem 2.13 for the duality of condensed Banach spaces and condensed Smith spaces in non-archimedean settings. For concreteness, we work over $\mathbb{Q}_{p}$ for some fixed prime $p$. As noted in [27, Lecture 6], the derived category $\mathsf{D(Solid}_{\mathbb{Q}_{p}})$ of solid $\mathbb{Q}_{p}$-vector spaces is a full subcategory of the derived category $\mathsf{D(Solid)}$ of solid abelian groups, which in turn is a full subcategory of the the derived category $\mathsf{D(Cond}(\mathsf{Ab}))$ of condensed abelian groups.
@@
As shown in [74, Lemma 3.8] (see also [27, Lecture 6], which more closely matches our notation), the solid $\mathbb{Q}_{p}$-Banach spaces are precisely the solid $\mathbb{Q}_{p}$-vector spaces of the form
@@
for some index set $I$ (readability dictates that for the remainder of this section we forego the underline notation for condensed images of rings). It is shown in [74, Lemma 3.10] that the classical Banach–Smith duality extends to the solid $\mathbb{Q}_{p}$-vector space setting via the internal $\mathrm{Hom}$ functor $\underline{\mathrm{Hom}}(-,-):(\mathsf{Solid}_{\mathbb{Q}_{p}})^{\mathrm{op}}\times\mathsf{Solid}_{\mathbb{Q}_{p}}\rightarrow\mathsf{Solid}_{\mathbb{Q}_{p}}$ (see [77, p. 13] for basic definitions of condensed internal $\mathrm{Hom}$ functors). In particular, there it is shown that, for any index set $I$, we have
@@
As $\prod_{I}\mathbb{Z}_{p}$ is, moreover, a projective solid $\mathbb{Z}_{p}$-module, it is essentially immediate that the second of these equalities holds even at the level of the derived category $\mathsf{D(Solid}_{\mathbb{Q}_{p}})$, i.e., even with the $\underline{\mathrm{Hom}}$ replaced by $\mathrm{R}\underline{\mathrm{Hom}}$; see [74, Remark 3.11]. The question of whether this holds for the first bullet (and hence of whether this is a derived duality relation), though, is more delicate. We end this subsection with a brief proof of the observation of Clausen and Scholze (again, see [27, Lecture 6]) that, under the conditions of Theorem 2.13 with $\lambda=\aleph_{0}$, this question also has an affirmative answer in the context of separable solid $\mathbb{Q}_{p}$-Banach spaces, i.e., in the situation in which the index set $I$ is countable. The following theorem yields clause (3) of Theorem B:
@@
By [74, Lemma 3.10], we have $\underline{\mathrm{Hom}}\left((\bigoplus_{\omega}\mathbb{Z}_{p})^{\wedge}_{p}\left[\frac{1}{p}\right],\mathbb{Q}_{p}\right)=(\prod_{\omega}\mathbb{Z}_{p})\left[\frac{1}{p}\right]$. Therefore, it suffices to show that $\mathrm{R}\underline{\mathrm{Hom}}\left((\bigoplus_{\omega}\mathbb{Z}_{p})^{\wedge}_{p}\left[\frac{1}{p}\right],\mathbb{Q}_{p}\right)$ is concentrated in degree zero. Note that
@@
For the second equality, see the third paragraph of [82, §4] (or, a little more concretely, argue as for [19, Lem. 11.1]). Since $S\in\mathsf{ED}$ and $\mathbb{Q}_{p}/(p^{m}\mathbb{Z}_{p})$ is discrete, we know that $\mathrm{Cont}(S,\mathbb{Q}_{p}/(p^{m}\mathbb{Z}_{p}))$ (with the compact-open topology) is a discrete abelian group for all $m<\omega$. Hence clause (3) of Theorem 2.13 applies to the interior RHom of line 2.4, entailing that
$\mathrm{H}^{i}\mathrm{RHom}\left(\lim_{n<\omega}\Big{(}\bigoplus_{\omega}\mathbb{Q}_{p}/(p^{n}\mathbb{Z}_{p})\Big{)},\mathrm{Cont}(S,\mathbb{Q}_{p}/(p^{m}\mathbb{Z}_{p}))\right)$